# Bayesian Component Separation for DESI LAE Automated Spectroscopic Redshifts and Photometric Targeting

### v1

@@ 4.2.1 Uniform noise
Using this procedure, we create 2,000 simulated LAE spectra and use our pipeline to automatically determine their redshifts using as and no sky component. An example simulated LAE spectrum can be seen in Figure 6. For each simulated spectrum, we inject a Ly emission line at a randomly selected redshift and calculate the signal-to-noise ratio (SNR) as:
@@
4.2.2 Injection into sky fibers
@@
For this, we randomly sample a redshift as in Equation 13 and inject the LAE signal scaled by a constant into a random sky residual spectrum from the selection outlined in Section 3.1, ensuring proper scaling using Equations 8 and 9 for inference in the transformed space. For this case, we calculate as the robust standard deviation (IQR/3.4896) of the sky residual spectrum between 3647 and 6078 Å, corresponding to the Ly wavelength at . Here, the scaled variance from the sky spectrum is used on the diagonal of the residual covariance .
@@
4.2.3 Error calibration
The and redshift uncertainty measurements are made under the assumption that the noise and uncertainties in the data are well modeled and a mean zero prior on each component. The model mismatch caused by assuming a mean zero prior for an emission line can distort the surface we measure as a function of redshift. In addition, there are many factors, such as stochasticity in the data as well as over- or under-estimation of the intrinsic noise, that could result in miscalibrated redshift errors from our pipeline. To correct for this surface distortion and calibrate the redshift errors we use our simulated LAE spectra created for injection tests to examine the calibration of our uncertainties and calculate a correction factor to be applied to the real data. For both kinds of injection tests, we generate 2,000 new simulated spectra with a higher SNR range (SNR ), and fit them with our pipeline as described above. For each simulated spectrum, the z-score was calculated as follows:
where is here the uncertainty on the redshift estimate as outlined in Section 3.3. Well-calibrated error bars yield z-scores that follow a unit normal distribution. To be robust to outliers, we evaluate this by calculating the interquartile range (IQR) of the distribution of z-scores and comparing it to the IQR of standard normal distribution (1.34896).
We calibrate our values by assuming that the surface distortion is a simple multiplicative correction, such that if the amplitude of the value is corrected, then the curvature will also be corrected. This allows us to derive a correction factor and independently verify that the resulting for the updated values is in fact well-described by a unit normal distribution. To correct the amplitude of the surface, we assume that the is dominated by removing the LAE signal which would have otherwise factored into the residuals. Then, the should be equal to the SNR, as defined in Section 4.2.1. We can compute the multiplicative factor by which this differ on injection tests, which we call , as:
@@ 4.2.3 Error calibration
For the constant-noise injection tests, we calculate , suggesting that the noise was well-modeled, and the adjusted z-score distribution yielded an IQR of 1.35/1.348 = 1.003. For the injection into real sky residual spectra, we calculate , and the adjusted IQR is 1.349/1.348 = 1.00. Both ratios are very close to 1, suggesting that after calibrating the surface, the adjusted redshift errors are also well-calibrated.
@@
We can use the factor calculated from the injection tests on real sky residual spectra to calibrate the reported and redshift uncertainties on the real LAE target spectra, using the VI redshifts as ground truth. For our dataset targeted with the Subaru Suprime-Cam, we apply the same corrections and recover an adjusted IQR of 2.00/1.348 = 1.48, where we exclude catastrophic outliers () from the IQR calculation to reduce bias.
This factor is used with Equations 20 and 21 to recalibrate the and redshift uncertainty for an LAE detection for the Subaru LAE candidates, as seen in Figures 4 and 5. A larger IQR ratio is expected for real data, which includes uncertainties associated with the VI redshifts as they are not ground truth. We can also fit an additional uncertainty term associated with the VI redshifts that is added in quadrature with the redshift uncertainties. Fitting this to an IQR ratio of 1 leads to an approximate VI redshift uncertainty of . This provides an approximate quantification of the potential uncertainties of the VI redshifts. Including this VI uncertainty in , we find that our pipeline determines redshifts within 3 of the VI redshift for 736 out of the 910 Subaru LAE targets ( 81%). This provides an additional metric of redshift accuracy of our pipeline, in addition to the previously reported accuracy of 90% of Subaru LAE targets being within 0.005 of the VI redshift.
4.3 Implications for DESI-2
@@
We first select a subset of all targets that were observed with 15-24 exposures and did not contain “LBG” in the VI comments, leaving 330 targets. Each target then has a set of at least 15 individual exposure spectra, their corresponding inverse variance values, and the length of each exposure. For exposures, we bootstrap (with replacement) exposures for each object, coadd them using inverse-variance weighting, and determine the redshift using our pipeline, using the sum of the inverse variances on the diagonal of the residual covariance. For each , we calculate the standard deviation of the redshift estimates and the average total exposure time (added across each subset of exposures) across 10 random sets of exposures.
@@
5 Discussion
5.1 Pros and cons of this analysis
@@
5.2 Pros and cons of MADGICS
@@
As with any numerical method, the derivation of uncertainties is critical to understanding its performance. In Sections 3.3 and 4.2 we discuss how we determine redshift uncertainties based on the curvature of the surface and validate them using injection tests. These uncertainties assume that the pipeline has determined the correct redshift and are not informative in the case of catastrophic failures. Additionally the MADGICS formalism encompasses not only posterior means for each component but also posterior covariances, which can be calculated as
@@ 5.3 Future Work
Our method can be applied to many astrophysical use cases outside of individual emission lines. MADGICS can fit multiple eigenvectors for each component, meaning it can encapsulate much more complex spectra, such as AGN spectra. A demonstration of the full expressivity of this method is forthcoming in future work. MADGICS can also be generalized for applications in photometry, with pixel-pixel covariance matrices for applications to star-galaxy separation and crowded-field photometry (Finkbeiner, 2025).
@@ Appendix A Woodbury Updates for Computational Speedup
In this way, we can calculate the without ever inverting a dense matrix ( is diagonal), which provides significant computational speedup. More details about the calculations within the MADGICS framework can be found in Saydjari et al. (2025, in prep).

### v2

@@ IV.2.1 Uniform noise
Using this procedure, we create 5,000 simulated LAE spectra and use our pipeline to automatically determine their redshifts using as and no sky component. An example simulated LAE spectrum can be seen in Figure 7. For each simulated spectrum, we inject a Ly emission line at a randomly selected redshift and calculate the signal-to-noise ratio (SNR) as:
@@
IV.2.2 Injection into sky fibers
@@
For this, we randomly sample a redshift as in Equation 13 and inject the LAE signal scaled by a constant into a random sky residual spectrum from the selection outlined in Section III.1, ensuring proper scaling using Equations 8 and 9 for inference in the transformed space. For this case, we calculate as the robust standard deviation (IQR/1.34896) of the sky residual spectrum between 3647 and 6078 Å, corresponding to the Ly wavelength at . Here, the scaled variance from the sky spectrum is used on the diagonal of the residual covariance .
@@
IV.2.3 Error calibration
The and redshift uncertainty measurements are made under the assumption that the noise and uncertainties in the data are well modeled and a mean zero prior on each component. The model mismatch caused by assuming a mean zero prior for an emission line can distort the surface we measure as a function of redshift. In addition, there are many factors, such as stochasticity in the data as well as over- or under-estimation of the intrinsic noise, that could result in miscalibrated redshift errors from our pipeline. To correct for this surface distortion and calibrate the redshift errors we use our simulated LAE spectra created for injection tests to examine the calibration of our uncertainties and calculate a correction factor to be applied to the real data. For both kinds of injection tests, we generate 5,000 new simulated spectra with a higher SNR range (SNR ), and fit them with our pipeline as described above. For each simulated spectrum, the z-score was calculated as follows:
where is here the uncertainty on the redshift estimate as outlined in Section III.3. Well-calibrated error bars yield z-scores that follow a unit normal distribution. To be robust to outliers, we evaluate this by calculating the interquartile range (IQR) of the distribution of z-scores and comparing it to the IQR of standard normal distribution (1.34896).
We calibrate our values by assuming that the surface distortion is a simple multiplicative correction, such that if the amplitude of the value is corrected, then the curvature will also be corrected. This allows us to derive a correction factor and independently verify that the resulting for the updated values is in fact well-described by a unit normal distribution. To correct the amplitude of the surface, we assume that the is dominated by removing the LAE signal which would have otherwise factored into the residuals. Then, the should be equal to the SNR, as defined in Section IV.2.1. We can compute the multiplicative factor by which this differ on injection tests, which we call , as:
@@ IV.2.3 Error calibration
For the constant-noise injection tests, we calculate , suggesting that the noise was well-modeled, and the adjusted z-score distribution yielded an IQR of 1.3276/1.34896 = 0.984. For the injection into real sky residual spectra, we calculate , and the adjusted IQR is 1.2426/1.34896 = 0.921. Both ratios are quite close to 1, suggesting that after calibrating the surface, the adjusted redshift errors are also well-calibrated.
@@
We can use the factor calculated from the injection tests on real sky residual spectra to calibrate the reported and redshift uncertainties on the real LAE target spectra, using the VI redshifts as ground truth. For our dataset targeted with the Subaru Suprime-Cam, we apply the same corrections and recover an adjusted IQR of 2.0214/1.34896 = 1.498, where we exclude catastrophic outliers () from the IQR calculation to reduce bias.
This factor is used with Equations 20 and 21 to recalibrate the and redshift uncertainty for an LAE detection for the Subaru LAE candidates, as seen in Figures 4 and 6. A larger IQR ratio is expected for real data, which includes uncertainties associated with the VI redshifts as they are not ground truth. We can also fit an additional uncertainty term associated with the VI redshifts that is added in quadrature with the redshift uncertainties. Fitting this to an IQR ratio of 1 leads to an approximate VI redshift uncertainty of . This provides an approximate quantification of the potential uncertainties of the VI redshifts. Including this VI uncertainty in , we find that our pipeline determines redshifts within 3 of the VI redshift for 720 out of the 881 Subaru LAE targets (81.7%). This provides an additional metric of redshift accuracy of our pipeline, in addition to the previously reported accuracy of 90% of Subaru LAE targets being within 0.005 of the VI redshift.
IV.3 Implications for DESI-2
@@
We first select a subset of all targets that were observed with 15-24 exposures and did not contain “LBG” in the VI comments, leaving 310 targets. Each target then has a set of at least 15 individual exposure spectra, their corresponding inverse variance values, and the length of each exposure. For exposures, we bootstrap (with replacement) exposures for each object, coadd them using inverse-variance weighting, and determine the redshift using our pipeline, using the inverted sum of the inverse variances on the diagonal of the residual covariance. For each , we calculate the standard deviation of the redshift estimates and the average total exposure time (added across each subset of exposures) across 10 random sets of exposures.
@@
V Discussion
V.1 Pros and cons of this analysis
@@
V.2 Pros and cons of MADGICS
@@
As with any numerical method, the derivation of uncertainties is critical to understanding its performance. In Sections III.3 and IV.2 we discuss how we determine redshift uncertainties based on the curvature of the surface and validate them using injection tests. These uncertainties assume that the pipeline has determined the correct redshift and are not informative in the case of catastrophic failures. Additionally the MADGICS formalism encompasses not only posterior means for each component but also posterior covariances, which can be calculated as
@@ V.3 Future Work
Our method can be applied to many astrophysical use cases outside of individual emission lines. MADGICS can fit multiple eigenvectors for each component, meaning it can encapsulate much more complex spectra, such as AGN spectra. A demonstration of the full expressivity of this method is forthcoming in future work. MADGICS can also be generalized for applications in photometry, with pixel-pixel covariance matrices for applications to star-galaxy separation and crowded-field photometry.
@@ VI Conclusions
The data for all figures in this paper are available at 10.5281/zenodo.15103552.
@@ Appendix A Woodbury Updates for Computational Speedup
In this way, we can calculate the without ever inverting a dense matrix ( is diagonal), which provides significant computational speedup.